OpenStudy (anonymous):

Someone please help me solve: Groupon negotiates coupons for items discounted 50-90% off by having thousands of subscribers. As a result, it experienced a dramatic increase in value from October 2010 to March 2011. The function G(x)=0 0.4x^2 + 1.8x + 6 Approximates the company’s value in billions of dollars x months after October 2010. (a) Evaluate G(0). Interpret the result. (b) Determine when Groupon’s value was about $15 billion.

OpenStudy (anonymous):

Can I get some assistance please?

OpenStudy (anonymous):

plug in a 0 for x.... G(0) = ???

OpenStudy (anonymous):

G(0) = 0.04(0)^2 + 1.8(0) + 6 =

OpenStudy (anonymous):

G(0)=0 0.4x^2 + 1.8(0) +6

OpenStudy (anonymous):

simplify.... what u got?

OpenStudy (anonymous):

i have no idea where to begin

OpenStudy (anonymous):

G=0+6

OpenStudy (anonymous):